OpenStudy (unklerhaukus):

\[\neg\left[(\forall \epsilon > 0)(\exists \delta > 0)(\forall x)[|x - a| < \delta \Rightarrow |f(x) - f(a)| < \epsilon]\right]\]

OpenStudy (unklerhaukus):

how to begin

OpenStudy (anonymous):

Ugh! Epsilon Delta....

terenzreignz (terenzreignz):

Well, slowly move the negation bar into the expression, except every time it passes a universal quantifier, change it to an existential,and every time it passes an existential quantifier, change it to universal :D

OpenStudy (unklerhaukus):

\[\neg\left[(\forall \epsilon > 0)(\exists \delta > 0)(\forall x)[|x - a| < \delta \Rightarrow |f(x) - f(a)| < \epsilon]\right]\]\[\qquad\qquad\downarrow\]\[(\exists \epsilon > 0)(\forall \delta > 0)(\exists x)\left[\neg[|x - a| < \delta \Rightarrow |f(x) - f(a)| < \epsilon]\right]\]

OpenStudy (unklerhaukus):

now how do i distribute the negation across the rest/

terenzreignz (terenzreignz):

Negation of something of the form p ⇒ q is p • ¬q

terenzreignz (terenzreignz):

in this case p is |x - a| < delta and q is |f(x) - f(a)| < epsilon

OpenStudy (unklerhaukus):

\[(\exists \epsilon > 0)(\forall \delta > 0)(\exists x)\left[\neg[|x - a| < \delta \Rightarrow |f(x) - f(a)| < \epsilon]\right]\]\[\downarrow\]\[(\exists \epsilon > 0)(\forall \delta > 0)(\exists x)[|x - a| < \delta \wedge\neg[ |f(x) - f(a)| < \epsilon]]\]

terenzreignz (terenzreignz):

Then, negation of < is ≥

OpenStudy (unklerhaukus):

\[\downarrow\]\[(\exists \epsilon > 0)(\forall \delta > 0)(\exists x)[|x - a| < \delta \wedge |f(x) - f(a)| ≥ \epsilon]\]

terenzreignz (terenzreignz):

That's it, I can't go further than that...

OpenStudy (unklerhaukus):

Thank you terenzreignz

terenzreignz (terenzreignz):

:D